The plates of a parallel – plate capacitor uses a disk with a radius of 10mm. The separation distance between the plates is 100 nm and between the plates is a sheet of mica. What is the capacitance of the capacitor? 


1
Expert's answer
2022-03-06T15:17:17-0500
"C=\frac{\epsilon\epsilon_0A}{d}"

"C=\frac{5*8.85*10^{-12}*3.14*0.005^2}{100*10^{-9}}\\
=3.5*10^{-8}\:\rm F=35\: nF"